# Experimental study of the distributions of off-diagonal scattering-matrix elements of quantum graphs with symplectic symmetry

### v1

@@ Abstract
We report on experimental studies of the distribution of the off-diagonal elements of the scattering ($S$) matrix of open microwave networks with symplectic symmetry and a chaotic wave dynamics. These consist of two geometrically identical subgraphs with unitary symmetry described by complex conjugate Hamiltonians, that are coupled by a pair of bonds. The results are compared to random-matrix theory (RMT) predictions obtained on the basis of the Heidelberg approach for the $S$ matrix of open quantum-chaotic systems. We demonstrate that deviations from RMT predictions observed in the distributions may be attributed to the fact that the subgraphs are not fully connected.
@@ I Introduction
Quantum graphs were introduced by Linus Pauling [84] as models for organic molecules. Due to their simplicity they are employed as basic mathematical objects [85, 86, 87, 88, 89, 90, 91, 92], also for physical networks in the limit where the lengths of the wires are much larger than their widths [63, 40] and have been used to simulate a large variety of systems, e.g., quantum circuits [93, 94], waveguide systems [95, 96, 97, 98, 99, 100] and realizations of high-dimensional multipartite quantum states [101]. To construct graphs with symplectic symmetry, we followed the procedure proposed in [43, 102]. It has been employed in [44, 82] and in [60] to investigate spectral properties and the distributions of the diagonal elements of the $S$ matrix, respectively, of such systems. To obtain information on the off-diagonal $S$-matrix elements of open quantum-chaotic systems with symplectic symmetry, we added scattering channels to the usual construction scheme.
@@
In Sec. II we review the experimental realization of quantum graphs belonging to the symplectic universality class. Then, in Sec. III we outline how we compute the eigenstates and $S$ matrix of the corresponding quantum graphs [104, 116]. In Sec. IV we present the experimental and numerical results for the distributions of the off-diagonal elements of the GSE graphs. We demonstrate that they do not agree with RMT predictions. This is attributed to the fact, that the two subgraphs that form the GSE graph are not fully connected. For a better understanding of the unexpected feature we in addition performed random-matrix simulations on the basis of the Heidelberg approach. Finally, we summarize our findings in Sec. VI.
II Experimental realization of GSE graphs
In the first experiments with microwave networks modeling quantum graphs with symplectic symmetry, referred to as GSE graphs in the following, put forward in Ref. [43] and in Refs. [44, 82], spectral properties were investigated and found to be well described by the GSE. The GSE graphs are designed such that the corresponding Hamiltonian [43, 44] attains in the basis $\mathcal{W}=\{|1\uparrow\rangle,|2\uparrow\rangle,\dots,|N\uparrow\rangle,|1\downarrow\rangle,|2\downarrow\rangle,\dots,|N\downarrow\rangle\}$ the form
@@ II Experimental realization of GSE graphs
The matrix $V$ couples the matrix $H_{0}$ of the spin-up system, which is drawn from the GUE, with its complex conjugate which describes the spin-down system. A Hamiltonian with this structure exhibits symplectic symmetry, $YH^{\text{T}}Y^{\text{T}}=H$ with $Y=-i\tau^{(2)}\otimes\mathds{1}_{N}$, where $\tau^{(2)}$ denotes the second Pauli matrix. It belongs to the GSE only for nonzero matrices $V$ [72], because only then it complies with the condition that in a system with symplectic symmetry spin-rotation symmetry is broken [72]. A scheme of the GSE graph used for the numerical and experimental studies is depicted in Fig. 1. It consists of two geometrically identical GUE subgraphs modeling $H_{0}$ and $H_{0}^{\ast}$ in Eq. (1), with the vertices marked by indices $i$ and $\bar{i}$, corresponding to spin orientations $\uparrow$ and $\downarrow$, respectively. Each of the subgraphs consists of $\mathcal{V}=9$ vertices and the whole graph has $\mathcal{B}=22$ bonds connecting them. The geometry of the graph is defined by the connectivity matrix $C$, which has vanishing diagonal elements $C_{\tilde{i}\tilde{i}}=0$ and its off-diagonal elements are $C_{\tilde{i}\tilde{j}}=1$ if vertices $\tilde{i}$ and $\tilde{j}$ are connected and $C_{\tilde{i}\tilde{j}}=0$ otherwise. Corresponding bond lengths coincide, $L_{ij}=L_{\bar{i}\bar{j}}$.
@@
The experimental graph is a microwave network, which is constructed from coaxial microwave cables that are connected by T joints, corresponding to the bonds and vertices, respectively. Time-reversal invariance violation is induced by T-shaped circulators of opposite orientation introduced at corresponding vertices, that lead to a directionality as illustrated by the blue arrows, and may be described by an additional phase $\pi/2$, respectively $-\pi/2$ in the waves. The subgraphs are connected at two vertex pairs marked by $i_{0}=2$, $\bar{j}_{0}=\bar{3}$ and $j_{0}=3$, $\bar{i}_{0}=\bar{2}$, respectively, that have same lengths $L_{i_{0}\bar{j}_{0}}=L_{j_{0}\bar{i}_{0}}$. The requirement $V=-V^{T}$, that is, $V_{i_{0}\bar{j}_{0}}=-V_{j_{0}\bar{i}_{0}}$, is realized with an additional phase of $\pi$ on one of the bonds, e.g., on that connecting $i_{0}$ and $\bar{j}_{0}$. Different realizations of the GSE graph were obtained by increasing the lengths of two corresponding bonds marked by circles in Fig. 1 with phase shifters by the same amount. The absorption strength was changed by introducing 1 dB to 30 dB attenuators.
@@
The experiments were conducted with four instead of two [60] antennas attached to four ports $P_{1},P_{2},P_{\bar{1}}$ and $P_{\bar{2}}$, respectively. For the measurements of the $S$-matrix elements the antennas were connected to an Agilent N5227A vector network analyzer (VNS) via SUCOFLEX126EA/11PC35/1PC35 coaxial-cables emitting microwaves into the resonator via one antenna $a$ and receiving it at an other one, $b$. The VNA measures relative phases and ratios of the outgoing and incoming microwaves as function of their frequency $f$, yielding the complex $S$ matrix elements $S_{ba}(f)$.
@@
Figure 1: Schematic view of the microwave network with symplectic symmetry constructed from coaxial cables that are connected at T joints. The blue arrows indicate the effect of the T-shaped circulators. The coaxial cables connecting subgraphs marked by $(3,\bar{2})$ and $(2,\bar{3})$ are connected to phase shifters, indicated by dashed-line bonds. Additional phase shifters are connected to the bonds marked by a circle for changing lengths of bonds. The four antennas are attached at ports $P_{1},P_{2},P_{\bar{1}}$ and $P_{\bar{2}}$.
@@
Figure 2: Density plot of the transmission amplitude $|S_{1\bar{1}}(f)|$ versus microwave frequency $f$ for fixed relative length differences $\Delta\tilde{l}$ between the bonds connecting vertices $2$ and $\bar{3}$, respectively, $3$ and $\bar{2}$.
@@ III Numerical analysis of the corresponding quantum graphs
Here, the coordinate $x$ varies along the bond from $x=0$ at vertex $\tilde{i}$ to $x=L_{\tilde{i}\tilde{j}}$ at vertex $\tilde{j}$, and $A=-A^{T}$ denotes the magnetic vector potential which induces violation of time–reversal invariance. Furthermore, $A_{ij}=-A_{\bar{i}\bar{j}}$ on corresponding bonds, since one of the GUE graphs should be the complex conjugate of the other one. On the bonds that couple the GUE graphs no magnetic potential is introduced, $A_{\tilde{i}\tilde{j}}=0$. The wave-function components are subject to boundary conditions imposed at the vertices that ensure continuity and conservation of the current [104]. The microwave networks model quantum graphs with Neumann boundary conditions at the vertices [40]. They constitute a special case of $\delta$-type boundary conditions [104, 117, 87, 88]. It should be noted that the wave equation for microwave networks is mathematically equivalent to that of the corresponding quantum graph only below the cutoff frequency for the first transverse electric mode, where only the fundamental transverse electromagnetic mode can propagate between the conductors [118, 119].
@@
Figure 3: Nearest-neighbor spacing distribution $P(s)$, integrated nearest-neighbor spacing distribution $I(s)$, number variance $\Sigma^{2}(L)$ and rigidity $\Delta_{3}(L)$, red with, blue without magnetic field. Black solid and dashed lines show the GSE and GUE curves, respectively. Turquoise lines show the results for the GSE graph with four coupling bonds instead of only two.
measures for the spectral properties are shown for the graph exhibited in Fig. 1, also for the case with no magnetic field, $A_{\tilde{i}\tilde{j}}=0\,\forall\tilde{i}\tilde{j}$. In both cases the Hamiltonian takes the form Eq. (1), however, $H_{0}=H_{0}^{\ast}$ in the latter. In addition, we performed numerical simulations for the GSE graph Fig. 1 with four coupling bonds instead of only two [102]. To obtain it, we removed the bonds between vertices $7$ and $8$ and between $\bar{7}$ and $\bar{8}$, respectively, and connected the vertex $7$ with the vertex $\bar{8}$ and $8$ with $\bar{7}$. The spectral properties of both graphs agree well, and for the short-range correlations also with the GSE curves. However for the long-range correlations clear deviations from the RMT predictions are observed. These are attributed to nonuniversal contributions. Namely, quantum graphs with Neumann boundary conditions comprise eigenfunctions that are localized only on a fraction of the bonds [76, 44].
@@
Figure 4: Top: Comparison of the experimental distributions of the real and imaginary parts for $S_{12}$ (black and red solid lines) and for $S_{1\bar{2}}$ (black and red dashed lines). Middle: Comparison of the distributions for the real and imaginary parts of $S_{12}$ (black and red lines) with the numerical results (turquoise line) obtained from Eq. (5) for the imaginary parts. The numerical distribution for the real part lies on top of the three other distributions and therefore it is not shown. Bottom: Same as top for $S_{1\bar{2}}$.
@@
Figure 5: Top: Comparison of the experimental distributions of the real and imaginary parts for $S_{11}$ (black and red) with the corresponding numerical ones (orange and turquoise) obtained from Eq. (5). Bottom: Comparison of the distributions for the real and imaginary parts of $S_{12}$ (black and red) with the numerical results (turquoise) obtained from Eq. (5). The numerical distribution for the real part is not shown, because it lies on top of the other three curves.
@@
Figure 6: Comparison of the distributions of the real part of $S_{12}$ (black) and $S_{1\bar{2}}$ (red) and the imaginary part of $S_{1\bar{2}}$ (turquoise) for four coupling cables obtained from the numerical computations for the quantum graph. The results for the distribution of the imaginary part of $S_{12}$ is not shown, because it lies on top of the other curves. Results are shown for absorption $\epsilon=0.0175$.
@@
Figure 7: Top: Comparison of the distributions of the real and imaginary parts for $S_{12}$ (solid and dashed lines) with two coupling bonds (black) with those for four coupling bonds (red) obtained from the numerical computations for the quantum graph with $\epsilon=0.0175$. Middle: Same as top for $S_{1\bar{2}}$. Bottom: Same as top for $S_{11}$.
@@
Figure 8: Top: Comparison of the experimental (black) and numerical (red) distributions of the modulus of $S_{11}$ with RMT simulations (blue). The values of the absorption strength induced in the experiments by attenuators, of $\epsilon$ in the quantum graphs and of the absorption parameter $\tau_{abs}$ in the RMT simulations are given in the panels. Bottom: Same as top for $S_{12}$.
@@
Figure 9: Top: Modulus of $S_{12}$ and $S_{1\bar{2}}$ (solid and dashed lines) for the cases with two coupling bonds (black) and four coupling bonds (red). For the case with four coupling cables the distributions are well described by RMT (turquoise histogram). Bottom: Comparison of the experimental distributions for the same experimental data as in Fig. 4 for the imaginary parts of $S_{12}$ and $S_{1\bar{2}}$ (black solid and dashed lines) with numerical results (green solid and dashed lines) and RMT simulations (red solid and dashed lines) obtained by inserting into the RMT $S$ matrix a GSE matrix of the form Eq. (1) where in $V$ all matrix elements were set to zero except the elements $(i_{0},\bar{j}),(\bar{i}_{0},j),(\bar{j},i_{0}),(j,\bar{i}_{0})$ for 5 values of $i_{0}$ and $j=1,\dots,N$ with $N=200$ in the $2N\times 2N$-dimensional Hamiltonian $\hat{H}$q. (1), for an ensemble of 300 $S$ matrices and $\tau_{\rm abs}=25$.

### v2

@@ Abstract
We report on experimental studies of the distribution of the off-diagonal elements of the scattering ($S$) matrix of open microwave networks with symplectic symmetry and a chaotic wave dynamics. These consist of two geometrically identical subgraphs with unitary symmetry described by complex conjugate Hamiltonians, that are coupled by a pair of bonds. The results are compared to random-matrix theory (RMT) predictions obtained on the basis of the Heidelberg approach for the $S$ matrix of open quantum-chaotic systems, employing random matrices from the Gaussian symplectic ensemble. We demonstrate that deviations observed in the distributions of the off-diagonal $S$-matrix elements may be attributed to the fact that the subgraphs are not fully connected, and propose a RMT model, which takes this into account and indeed confirms the experimental results.
@@ I Introduction
Quantum graphs were introduced by Linus Pauling [84] as models for organic molecules. Due to their simplicity they are employed as basic mathematical objects [85, 86, 87, 88, 89, 90, 91, 92], also for physical networks in the limit where the lengths of the wires are much larger than their widths [63, 40] and have been used to simulate a large variety of systems, e.g., quantum circuits [93, 94], waveguide systems [95, 96, 97, 98, 99, 100] and realizations of high-dimensional multipartite quantum states [101]. Following the procedure proposed in [43, 102], we construct graphs with symplectic symmetry using two geometrically identical subgraphs that are coupled by a pair of bonds. It has been employed in [44, 82] and in [60] to investigate spectral properties and the distributions of the diagonal elements of the $S$ matrix, respectively, of such systems. To obtain information on the off-diagonal $S$-matrix elements of open quantum-chaotic systems with symplectic symmetry, we added scattering channels to the usual construction scheme.
@@
In Sec. II we review the experimental realization of quantum graphs with symplectic symmetry, referred to as symplectic graphs in the following. Then, in Sec. III we outline how we compute the eigenstates and $S$ matrix of the corresponding quantum graphs [104, 116]. In Sec. IV we present the experimental and numerical results for the distributions of the off-diagonal elements of the $S$ matrix of these symplectic graphs. We demonstrate that they do not agree with RMT predictions deduced from the Heidelberg approach, which employs Hermitean random matrices from the GSE. This is attributed to the fact, that the two subgraphs that form the symplectic graph are not fully connected. To confirm this assumption, we in addition performed simulations with random $S$ matrices constructed on the basis of the same scattering formalism, however with modified Hermitean random matrices, that resemble the symplectic graph. Finally, we summarize our findings in Sec. VI.
II Experimental realization of graphs with symplectic symmetry
In the first experiments with microwave networks modeling quantum graphs with symplectic symmetry, put forward in Ref. [43] and in Refs. [44, 82], spectral properties were investigated and found to be well described by the GSE. The symplectic graphs are designed such that the corresponding Hamiltonian [43, 44] attains in the basis $\mathcal{W}=\{|1\uparrow\rangle,|2\uparrow\rangle,\dots,|N\uparrow\rangle,|1\downarrow\rangle,|2\downarrow\rangle,\dots,|N\downarrow\rangle\}$ the form
@@ II Experimental realization of graphs with symplectic symmetry
The matrix $V$ couples the matrix $H_{0}$ of the spin-up system, which is drawn from the GUE, with its complex conjugate which describes the spin-down system. A Hamiltonian with this structure exhibits symplectic symmetry, $YH^{\text{T}}Y^{\text{T}}=H$ with $Y=-i\tau^{(2)}\otimes\mathds{1}_{N}$, where $\tau^{(2)}$ denotes the second Pauli matrix. It belongs to the GSE only for nonzero matrices $V$ [72], because only then it complies with the condition that in a system with symplectic symmetry spin-rotation symmetry is broken [72].
A scheme of the symplectic graph used for the numerical and experimental studies is depicted in Fig. 1. It consists of two geometrically identical GUE subgraphs belonging to the unitary universality class and modeling $H_{0}$ and $H_{0}^{\ast}$ in Eq. (1), with the vertices marked by indices $i$ and $\bar{i}$, corresponding to spin orientations $\uparrow$ and $\downarrow$, respectively. Each of the subgraphs consists of $\mathcal{V}=9$ vertices and the whole graph has $\mathcal{B}=22$ bonds connecting them. The geometry of the graph is defined by the connectivity matrix $C$, which has vanishing diagonal elements $C_{\tilde{i}\tilde{i}}=0$ and its off-diagonal elements are $C_{\tilde{i}\tilde{j}}=1$ if vertices $\tilde{i}$ and $\tilde{j}$ are connected and $C_{\tilde{i}\tilde{j}}=0$ otherwise. Corresponding bond lengths coincide, $L_{ij}=L_{\bar{i}\bar{j}}$.
@@
The experimental graph is a microwave network, which is constructed from coaxial microwave cables that are connected by T joints, corresponding to the bonds and vertices, respectively. Time-reversal invariance violation is induced by T-shaped circulators of opposite orientation introduced at corresponding vertices, that lead to a directionality as illustrated by the blue arrows, and may be described by an additional phase $\pi/2$, respectively $-\pi/2$ in the waves. The subgraphs are connected at two vertex pairs marked by $i_{0}=2$, $\bar{j}_{0}=\bar{3}$ and $j_{0}=3$, $\bar{i}_{0}=\bar{2}$, respectively, that have same lengths $L_{i_{0}\bar{j}_{0}}=L_{j_{0}\bar{i}_{0}}$. The requirement $V=-V^{T}$, that is, $V_{i_{0}\bar{j}_{0}}=-V_{j_{0}\bar{i}_{0}}$, is realized with an additional phase of $\pi$ on one of the bonds, e.g., on that connecting $i_{0}$ and $\bar{j}_{0}$. Different realizations of the symplectic graph were obtained by increasing the lengths of two corresponding bonds marked by circles in Fig. 1 with phase shifters by the same amount. The absorption strength was changed by introducing 1 dB to 30 dB attenuators.
@@
The experiments were conducted with four instead of two [60] antennas attached to four ports $P_{1},P_{2},P_{\bar{1}}$ and $P_{\bar{2}}$, respectively. For the measurements of the $S$-matrix elements the antennas were connected to an Agilent N5227A vector network analyzer (VNS) via SUCOFLEX126EA/11PC35/1PC35 coaxial-cables emitting microwaves into the resonator via one antenna $a$ and receiving it at an other one, $b$. The VNA measures relative phases and ratios of the outgoing and incoming microwaves as function of their frequency $f$, yielding the complex $S$-matrix elements $S_{ba}(f)$.
@@
Figure 1: Schematic view of the microwave network with symplectic symmetry constructed from coaxial cables that are connected at T joints. The blue arrows indicate the effect of the T-shaped circulators. The coaxial cables connecting subgraphs marked by $(3,\bar{2})$ and $(2,\bar{3})$ are connected to phase shifters, indicated by dashed-line bonds. Additional phase shifters are connected to the bonds marked by circles to change the lengths of bonds. The four antennas are attached at ports $P_{1},P_{2},P_{\bar{1}}$ and $P_{\bar{2}}$.
@@
Figure 2: Density plot of the transmission amplitude $|S_{1\bar{1}}(f)|$ versus microwave frequency $f$ for fixed relative length differences $\Delta\tilde{l}$ between the bonds connecting vertices 222 and $\bar{3}$, respectively, 333 and $\bar{2}$.
@@ III Numerical analysis of the corresponding quantum graphs
Here, the coordinate $x$ varies along the bond from $x=0$ at vertex $\tilde{i}$ to $x=L_{\tilde{i}\tilde{j}}$ at vertex $\tilde{j}$, and $A=-A^{T}$ denotes the magnetic vector potential which induces violation of time–reversal invariance. Furthermore, $A_{ij}=-A_{\bar{i}\bar{j}}$ on corresponding bonds, since one of the GUE graphs should be the complex conjugate of the other one. On the bonds that couple the GUE graphs we set $A_{\tilde{i}\tilde{j}}=0$. The wave-function components are subject to boundary conditions imposed at the vertices that ensure continuity and conservation of the current [104]. The microwave networks model quantum graphs with Neumann boundary conditions at the vertices [40]. They constitute a special case of $\delta$-type boundary conditions [104, 117, 87, 88]. It should be noted that the wave equation for microwave networks is mathematically equivalent to that of the corresponding quantum graph only below the cutoff frequency for the first transverse electric mode, where only the fundamental transverse electromagnetic mode can propagate between the conductors [118, 119].
@@
Figure 3: Nearest-neighbor spacing distribution $P(s)$, integrated nearest-neighbor spacing distribution $I(s)$, number variance $\Sigma^{2}(L)$ and rigidity $\Delta_{3}(L)$, with (red lines) and without (turquoise lines with triangles) magnetic field. Black solid and dashed lines show the GSE and GUE curves, respectively. Blue lines with squares show the results for the symplectic graph with four coupling bonds instead of only two.
measures for the spectral properties are shown for the graph exhibited in Fig. 1, also for the case with no magnetic field, $A_{\tilde{i}\tilde{j}}=0\,\forall\tilde{i}\tilde{j}$. In both cases the Hamiltonian takes the form Eq. (1), however, $H_{0}=H_{0}^{\ast}$ in the latter case. In addition, we performed numerical simulations for the GUE graph Fig. 1 with four coupling bonds instead of only two [102]. To obtain it, we removed the bonds between vertices 777 and 888 and between $\bar{7}$ and $\bar{8}$, respectively, and connected the vertex 777 with the vertex $\bar{8}$ and 888 with $\bar{7}$. The spectral properties of both graphs agree well, for the short-range correlations also with the GSE curves, that is, the level repulsion is quartic, thus corroborating that the graph exhibits symplectic symmetry. Furthermore, this indicates high spectral rigidity which is reflected in the logarithmic dependence of the number variance $\Sigma^{2}(L)$ and the rigidity $\Delta_{3}(L)$ on the interval length $L$, that is, number of mean spacings, in contrast to a linear one exhibited by uncorrelated levels [120, 121]. These are shown in the lower part of Fig. 3. Clear deviations of the long-range correlations from the RMT predictions are observed beyond $L\gtrsim 2.5$ for $\Sigma^{2}(L)$ and $L\gtrsim 5$ for $\Delta_{3}(L)$. They are attributed to nonuniversal contributions. Namely, quantum graphs with Neumann boundary conditions comprise eigenfunctions that are localized only on a fraction of the bonds [76, 44].
@@
Figure 4: Top: Comparison of the experimental distributions of the real and imaginary parts for $S_{12}$ (black line and red solid line with dots) and for $S_{1\bar{2}}$ (black dashed line with crosses and red dashed line with dots). Middle: Comparison of the distributions for the real and imaginary parts of $S_{12}$ (black line and red line with dots) with the numerical results (turquoise line with triangles) for the corresponding quantum graph obtained from Eq. (5) for the imaginary parts. The numerical distribution for the real part lies on top of the three other distributions and therefore it is not shown. Bottom: Same as top for $S_{1\bar{2}}$.
@@
Figure 5: Top: Comparison of the experimental distributions of the real and imaginary parts for $S_{11}$ (black line and red line with dots) with the corresponding numerical ones for the corresponding quantum graph (orange line with triangles down and turquoise line with triangles up) obtained from Eq. (5). Bottom: Comparison of the distributions for the real and imaginary parts of $S_{12}$ (black line and red line with dots) with the numerical results (turquoise line with triangles) obtained from Eq. (5). The numerical distribution for the real part is not shown, because it lies on top of the other three curves.
@@
Figure 6: Comparison of the distributions of the real part of $S_{12}$ (black line) and $S_{1\bar{2}}$ (red line with dots) and the imaginary part of $S_{1\bar{2}}$ (turquoise line with triangles) for four coupling cables obtained from the numerical computations for the quantum graph. The results for the distribution of the imaginary part of $S_{12}$ is not shown, because it lies on top of the other curves. Results are shown for absorption $\epsilon=0.0175$.
@@
Figure 7: Top: Comparison of the distributions of the real and imaginary parts for $S_{12}$ (solid and dashed lines) with two coupling bonds (black solid line and dashed line with crosses) with those for four coupling bonds (red lines with dots) obtained from the numerical computations for the quantum graph with $\epsilon=0.0175$. Middle: Same as top for $S_{1\bar{2}}$. Bottom: Same as top for $S_{11}$.
@@
Figure 8: Top: Comparison of the experimental (black lines) and numerical (red lines with dots) distributions of the modulus of $S_{11}$ with RMT simulations (blue lines with crosses). The values of the absorption strength induced in the experiments by attenuators, of $\epsilon$ in the quantum graphs and of the absorption parameter $\tau_{abs}$ in the RMT simulations are given in the panels. Bottom: Same as top for $S_{12}$.
@@
Figure 9: Top: Modulus of $S_{12}$ and $S_{1\bar{2}}$ (solid lines and dashed lines) for the cases with two coupling bonds (black solid line and dashed line with crosses) and four coupling bonds (red lines with dots). For the case with four coupling cables the distributions are well described by RMT (turquoise line with triangles). Bottom: Comparison of the experimental distributions for the same experimental data as in Fig. 4 for the imaginary parts of $S_{12}$ and $S_{1\bar{2}}$ (black solid line and dashed line with crosses) with numerical results for the corresponding quantum graph (green solid and dashed lines with triangles) and RMT simulations (red solid and dashed lines with dots) obtained by inserting into the RMT $S$ matrix a GSE matrix of the form Eq. (1) where in $V$ all matrix elements were set to zero except the elements $(i_{0},\bar{j}),(\bar{i}_{0},j),(\bar{j},i_{0}),(j,\bar{i}_{0})$ for 5 values of $i_{0}$ and $j=1,\dots,N$ with $N=200$ in the $2N\times 2N$-dimensional Hamiltonian $\hat{H}$, Eq. (1), for an ensemble of 300 $S$ matrices and $\tau_{\rm abs}=25$.